# Multimodal Prompt Optimization: Why Not Leverage Multiple Modalities for MLLMs

### v1

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated outstanding capabilities across a diverse range of tasks and domains (OpenAI, 2024; Grattafiori et al., 2024; Yang et al., 2025). We note that a central factor in unlocking their full potential lies in the design of prompts, which directly influence model performance. However, crafting high-quality prompts is often a labor-intensive and iterative process that requires substantial human intervention. To address this limitation, the field of Automatic Prompt Optimization (APO) has emerged, whose goal is to automate the discovery of effective prompts (Zhou et al., 2023; Pryzant et al., 2023; Yang et al., 2024; Fernando et al., 2024). For example, one representative approach (APE) frames this challenge as an iterative search problem, where at each step, a set of new candidate prompts is generated or updated, evaluated on a target task, and the best-performing prompts are selected to guide the next round of generation.
Recently, on top of the LLMs, Multimodal Large Language Models (MLLMs) have been proposed, which process not only text but also images, videos, and other modalities (such as molecules) (OpenAI, 2023; Zhu et al., 2025; Bai et al., 2025; Gemini, 2025). Yet, despite these advances and their wide-ranging applications, existing prompt optimization methods remain restricted to the textual modality (Pryzant et al., 2023; Guo et al., 2024; Cui et al., 2025), and overlook the richer expressive capacity afforded by multimodal inputs (that text alone cannot capture). For instance, as illustrated in Figure 1, describing the distinct characteristics of a specific bird may require long and potentially ambiguous text, while a single image can convey the same information far more directly. By limiting optimization to text, existing methods are prone to generating less effective, suboptimal prompts that fail to fully exploit the multimodal space that MLLMs are inherently capable of leveraging.
Motivated by this limitation, we first define the novel problem of multimodal prompt optimization, which expands the prompt optimization space beyond text to incorporate multiple modalities. However, while this expanded space opens new opportunities, it also introduces a couple of challenges for automatic optimization. First, exploring the larger, combinatorial space of multimodal prompts requires a prompt update strategy that can efficiently navigate candidate prompts while maintaining cross-modal consistency. Furthermore, selecting promising candidates becomes substantially more difficult, as the enlarged search space makes optimal prompts increasingly sparse, given the need to account for both the effectiveness within each modality and the alignment across modalities, which, in turn, calls for evaluation strategies that are both efficient and accurate.
@@ 2 Related Work
The development of MLLMs has significantly extended the capabilities of traditional LLMs by enabling them to process and reason over diverse non-textual modalities, including images, videos, audio, and more (Liu et al., 2023; OpenAI, 2023; Chu et al., 2023). In particular, these models are typically trained through large-scale multimodal pre-training, which aligns modality-specific encoders (e.g., vision or audio) with LLM backbones, followed by post-training stages such as supervised fine-tuning and preference optimization to endow them with multimodal instruction-following abilities (Gemini, 2025; Bai et al., 2025; Zhu et al., 2025). Moreover, leveraging these capabilities, MLLMs have achieved strong performance on a broad range of tasks, from foundational ones such as classification and captioning, to domain-specific, high-stakes applications such as medical image question answering and pharmacological property prediction (Martin et al., 2019; Liu et al., 2021; Corbière et al., 2025; Huang et al., 2021).
@@
To reduce the burden of manual prompt engineering and systematically uncover the effective prompts, the field of Automatic Prompt Optimization (APO) has emerged. Existing works can be broadly categorized into two paradigms. The first is gradient-based optimization, which learns continuous embedding vectors (i.e., soft prompts) that are prepended to model inputs to steer behavior (Khattak et al., 2023; Zeng et al., 2024; Wang et al., 2024). Yet, while effective, they are computationally costly, yield uninterpretable numerical vectors, and are restricted to open-source models with accessible parameters. To overcome these drawbacks, gradient-free approaches have been proposed, which iteratively generate, evaluate, and refine candidate prompts using LLMs themselves (Zhou et al., 2023; Yang et al., 2024). Also, some recent works enhance this process by analyzing prompt failures to guide improvements (Pryzant et al., 2023; Ye et al., 2024; Cui et al., 2025; Yuksekgonul et al., 2025), while others borrow ideas from evolutionary algorithms (e.g., mutation and crossover) to explore the prompt space (Guo et al., 2024; Fernando et al., 2024). Despite this progress, current APO techniques are limited to text-only settings, restricting optimization to purely linguistic information. In contrast, our work expands prompt optimization into the multimodal domain, enabling the prompt discovery that fully exploits the capabilities of MLLMs.
@@ 3.3 Effective Prompt Selection by Prior-Inherited Bayesian UCB
Another challenge in multimodal prompt optimization is to identify which candidates should be prioritized for evaluation and carried forward. Yet, this step is non-trivial with the enlarged multimodal space, since high-quality prompts become relatively sparse, and a large portion of the evaluation budget risks being wasted on low-potential candidates. Existing approaches typically adopt either (i) uniform allocation, where each candidate is evaluated equally regardless of its prior likelihood of success (Zhou et al., 2023; Cui et al., 2025), or (ii) bandit-based allocation, such as UCB (Auer, 2002; Pryzant et al., 2023), which adaptively balances exploration and exploitation. However, both paradigms suffer from an inefficient cold-start problem: newly generated prompts are treated as independent arms with no prior information, leading to unproductive evaluations in the early rounds.
@@
Motivated by this finding, we propose prior-inherited Bayesian UCB, a selection strategy that initializes the score distribution of a new child prompt based on the posterior of its parent (rather than uniform). Specifically, we model the expected score of each multimodal prompt as a Beta distribution, , where and correspond to (pseudo-) counts of successful and failure outcomes, respectively. Then, for a child prompt originated from a parent prompt , we initialize its prior proportionally to the posterior mean performance of the parent , scaled by a prior strength hyperparameter , formalized as follows:
@@ 4 Experiments
4.1 Experimental Setup
Datasets
Baselines
We benchmark MPO against both manually designed prompts and representative automatic prompt optimization methods. For manual prompting, we include Human, a simple hand-crafted prompt, Chain-of-Thought (CoT) (Wei et al., 2022), which uses the widely adopted phrase “Let’s think step by step,” and Few-Shot, which supplies in-context examples drawn from the training data. For automatic methods, we compare against leading LLM-based text-only optimizers, including APE (Zhou et al., 2023), OPRO (Yang et al., 2024), EvoPrompt (Guo et al., 2024), PE2 (Ye et al., 2024), ProTeGi (Pryzant et al., 2023), and SEE (Cui et al., 2025). Detailed descriptions of all baselines are provided in Appendix A.2.
Implementation Details
@@ 4.2 Experimental Results and Analyses
Recall that MPO uses the alignment-preserving exploration to jointly refine textual and non-textual components of multimodal prompts, and we further analyze how this cross-modal alignment strategy contributes to performance gains. To isolate this effect, we consider four variants: (1) Sequential, where the textual prompt is optimized first and the non-textual prompt is refined afterward; (2) Random Image Prompt, where the image component is replaced with another optimized image prompt (i.e., not jointly optimized with the text); (3) In-Distribution Image Query, where it is replaced with an image sampled from the same task; and (4) OOD Image Query, where it is replaced with an image sampled from a different task. After that, we measure the relationship between performance gain over the Human baseline and the DSG score designed to quantify alignment (Cho et al., 2024). As shown in Figure 4, MPO achieves both the highest alignment score and the largest performance gains, followed by Sequential optimization and Random Image Prompt, while In-Distribution and OOD Image Query lag significantly behind. These results confirm that stronger cross-modal alignment directly translates to better task performance, and that alignment-preserving updates (included in MPO) are crucial in promoting modality consistency.
Ablation on Modality Contributions in Prompts
@@ 5 Conclusion
We attach the code to ensure the reproducibility of our work in the supplementary materials. Additionally, we provide a detailed description of the experimental setup in Section 4.1. We further provide additional implementation details in Appendix A.3, the dataset configuration in Appendix A.1, the meta prompts to operationalize MPO in Appendix A.4, and the full algorithms in Appendix A.5.
@@ A.2 Details on Baselines
SEE (Cui et al., 2025) performs cohesive optimization of both the prompt instructions and the in-context examples. The method follows a four-phase process that strategically alternates between global exploration and local exploitation.
@@ C.3 Qualitative Results
In this section, we provide qualitative results for MPO. Figure 14 illustrates the optimization process in the molecular domain, and Table 8 shows examples of textual conditions for the modality-specific generator and the resulting image prompts. The optimized multimodal prompts from MPO are presented for the image (Table 9), video (Table 10), and molecular domains (Table 11 and Table 12).

### v2

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated outstanding capabilities across a diverse range of tasks and domains (OpenAI, 2024; Grattafiori and others, 2024; Yang et al., 2025). We note that a central factor in unlocking their full potential lies in the design of prompts, which directly influence model performance. However, crafting high-quality prompts is often a labor-intensive and iterative process that requires substantial human intervention. To address this limitation, the field of Automatic Prompt Optimization (APO) has emerged, whose goal is to automate the discovery of effective prompts (Pryzant et al., 2023; Ramnath et al., 2025; Cui et al., 2025a; Li et al., 2025). For example, Zhou et al. (2023) frames this challenge as an iterative search problem, where at each step, a set of new candidate prompts is generated or updated, evaluated on a target task, and the best-performing prompts are selected to guide the next round of generation.
Recently, on top of the LLMs, Multimodal Large Language Models (MLLMs) have been proposed, which process not only text but also images, videos, and other modalities (such as molecules) (Kim et al., 2025; Zheng et al., 2024; Song et al., 2025). Yet, despite these advances and their wide-ranging applications, existing prompt optimization methods remain restricted to the textual modality (Choi et al., 2025; Fernando et al., 2024; Cui et al., 2025b), and overlook the richer expressive capacity afforded by multimodal inputs (that text alone cannot capture). For instance, as illustrated in Figure 1, describing the distinct characteristics of a specific bird may require long and potentially ambiguous text, while a single image can convey the same information far more directly. By limiting optimization to text, existing methods are prone to generating less effective, suboptimal prompts that fail to fully exploit the multimodal space that MLLMs are inherently capable of leveraging.
Motivated by this limitation, we first define the novel problem of multimodal prompt optimization111We define a multimodal prompt as a pair of a discrete textual and non-textual prompts., which expands the prompt optimization space beyond text to incorporate multiple modalities. However, while this expanded space opens new opportunities, it also introduces a couple of challenges for automatic optimization. First, exploring the larger, combinatorial space of multimodal prompts requires a prompt update strategy that can efficiently navigate candidate prompts while maintaining cross-modal consistency. Furthermore, selecting promising candidates becomes substantially more difficult, as the enlarged search space makes optimal prompts increasingly sparse, given the need to account for both the effectiveness within each modality and the alignment across modalities, which, in turn, calls for evaluation strategies that are both efficient and accurate.
@@ 2 Related Work
The development of MLLMs has significantly extended the capabilities of traditional LLMs by enabling them to process and reason over diverse non-textual modalities, including images, videos, audio, and more (Liu et al., 2023; Chu et al., 2023; M. Bran et al., 2024). In particular, these models are typically trained through large-scale multimodal pre-training, which aligns modality-specific encoders (e.g., vision or audio) with LLM backbones, followed by post-training stages such as supervised fine-tuning and preference optimization to endow them with multimodal instruction-following abilities (Gemini, 2025; Bai et al., 2025; Zhu et al., 2025). Moreover, leveraging these capabilities, MLLMs have achieved strong performance on a broad range of tasks, from foundational ones such as classification and captioning, to domain-specific, high-stakes applications such as medical image question answering and pharmacological property prediction (Martin et al., 2019; Liu et al., 2021; Corbière et al., 2025; Huang et al., 2021).
@@
To reduce the burden of manual prompt engineering and systematically uncover the effective prompts, the field of Automatic Prompt Optimization (APO) has emerged. Existing works can be broadly categorized into two paradigms. The first is gradient-based optimization, which learns continuous embedding vectors (i.e., soft prompts) that are prepended to model inputs to steer behavior (Khattak et al., 2023; Zeng et al., 2024; Wang et al., 2024). Yet, while effective, they are computationally costly, yield uninterpretable numerical vectors, and are restricted to open-source models with accessible parameters. To overcome these drawbacks, gradient-free approaches have been proposed, which iteratively generate, evaluate, and refine candidate prompts using LLMs themselves (Zhou et al., 2023; Yang et al., 2024). Also, some recent works enhance this process by analyzing prompt failures to guide improvements (Khattab et al., 2024; Ye et al., 2024; Cui et al., 2025b; Yuksekgonul et al., 2025), while others borrow ideas from evolutionary algorithms (e.g., mutation and crossover) to explore the prompt space (Guo et al., 2024; Fernando et al., 2024). Despite this progress, current APO techniques are limited to text-only settings, restricting optimization to purely linguistic information. In contrast, our work expands prompt optimization into the multimodal domain, enabling the prompt discovery that fully exploits the capabilities of MLLMs.
@@ 3.2 Alignment-Preserving Exploration of Multimodal Prompt Space
For the optimization process, we adopt the beam search. Specifically, at each iteration, we select top- best-performing multimdoal prompts , and then apply the cohesive backpropagation and the joint multimodal update to generate new multimodal prompts from the top- multimodal prompts . We refer to and as parent and child prompts, respectively.
@@
It is worth noting that, for the generation and edit operators, we randomly select one parent prompt from top- multimodal prompts of the previous iteration to generate a child prompt (i.e., , while parent prompts are selected for the mix operator (i.e., ).
@@ 3.3 Effective Prompt Selection by Prior-Inherited Bayesian UCB
Another challenge in multimodal prompt optimization is to identify which candidates should be prioritized for evaluation and carried forward (Pryzant et al., 2023; Shi et al., 2024). Yet, this step is non-trivial with the enlarged multimodal space, since high-quality prompts become relatively sparse, and a large portion of the evaluation budget risks being wasted on low-potential candidates. Existing approaches typically adopt either (i) uniform allocation, where each candidate is evaluated equally regardless of its prior likelihood of success (Zhou et al., 2023; Cui et al., 2025b), or (ii) bandit-based allocation, such as UCB (Auer, 2002; Bouneffouf and Féraud, 2025; Li et al., 2026; Ashizawa et al., 2025), which adaptively balances exploration and exploitation. However, both paradigms suffer from an inefficient cold-start problem: newly generated prompts are treated as independent arms with no prior information, leading to unproductive evaluations in the early rounds.
@@
Motivated by this finding, we propose prior-inherited Bayesian UCB (Kaufmann et al., 2012), a selection strategy that initializes the score distribution of a new child prompt based on the posterior of its parent (rather than uniform). Specifically, we model the expected score of each multimodal prompt as a Beta distribution, , where and correspond to (pseudo-) counts of successful and failure outcomes, respectively. Then, for a child prompt originated from a parent prompt 222If a child prompt is generated by the mix operator, we average the posterior mean performance of multiple parent prompts., we initialize its prior proportionally to the posterior mean performance of the parent , scaled by a prior strength hyperparameter , formalized as follows:
@@ 4 Experiments
4.1 Experimental Setup
Datasets
Baselines
We benchmark MPO against both manually designed prompts and representative automatic prompt optimization methods. For manual prompting, we include Human, a simple hand-crafted prompt, Chain-of-Thought (CoT) (Wei et al., 2022), which uses the widely adopted phrase “Let’s think step by step,” and Few-Shot, which supplies in-context examples drawn from the training data. For automatic methods, we compare against leading LLM-based text-only optimizers, including APE (Zhou et al., 2023), OPRO (Yang et al., 2024), EvoPrompt (Guo et al., 2024), PE2 (Ye et al., 2024), ProTeGi (Pryzant et al., 2023), and SEE (Cui et al., 2025b). Detailed descriptions of all baselines are provided in Appendix A.2.
Implementation Details
@@ 4.2 Experimental Results and Analyses
Recall that MPO uses the alignment-preserving exploration to jointly refine textual and non-textual components of multimodal prompts, and we further analyze how this cross-modal alignment strategy contributes to performance gains. To isolate this effect, we consider four variants: (1) Sequential, where the textual prompt is optimized first and the non-textual prompt is refined afterward; (2) Random Image Prompt, where the image component is replaced with another optimized image prompt (i.e., not jointly optimized with the text); (3) In-Distribution Image Query, where it is replaced with an image sampled from the same task; and (4) OOD Image Query, where it is replaced with an image sampled from a different task. After that, we measure the relationship between performance gain over the Human baseline, as well as the DSG score333DSG decomposes the textual description into atomic, dependency-aware queries and verifies each using an MLLM, enabling a precise assessment of whether the visual content reflects the intended textual details. (Cho et al., 2024), used as a standard metric for measuring cross-modal alignment following prior work (Mañas et al., 2024). As shown in Figure 4, MPO achieves both the highest alignment score and the largest performance gains, followed by Sequential optimization and Random Image Prompt, while In-Distribution and OOD Image Query lag significantly behind. These results coAnfirm that stronger cross-modal alignment directly translates to better task performance, and that alignment-preserving updates (included in MPO) are crucial in promoting modality consistency.
@@ 5 Conclusion
We provide the public GitHub link to the code to ensure the reproducibility of our work. Additionally, we provide a detailed description of the experimental setup in Section 4.1. We further provide additional implementation details in Appendix A.3, the dataset configuration in Appendix A.1, the meta prompts to operationalize MPO in Appendix A.4, and the full algorithms in Appendix A.5.
Acknowledgement
This work was supported by the Institute for Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Program (KAIST) and RS-2022-II220713, Meta-learning Applicable to Real-world Problems), the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2023-00256259), the Korea Machine Learning Ledger Orchestration for Drug Discovery Project (K-MELLODDY) funded by the Ministry of Health & Welfare and Ministry of Science and ICT, Republic of Korea (No. RS-2024-00460870), the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (RS-2024-00414751), and the InnoCORE program of the Ministry of Science and ICT (No. N10250156).
@@ A.2 Details on Baselines
SEE (Cui et al., 2025b) performs cohesive optimization of both the prompt instructions and the in-context examples. The method follows a four-phase process that strategically alternates between global exploration and local exploitation.
@@ C.3 Qualitative Results
In this section, we present an additional comprehensive qualitative analysis of MPO.
Additional Analysis in MPO
Examples of Optimized Multimodal Prompts
Comparison with Text-only Optimization Baseline
We compare the optimized prompts and responses of MPO with those of a leading text-only optimization baseline (SEE) to investigate the underlying reasons for the effectiveness of using multimodal prompts. As shown in Tables 13 and 15, text-only methods must encode all necessary visual cues solely through language, often resulting in verbose descriptions and overly generic instructions (e.g., “Pay attention to specific visual markers such as bill shape, color, and the feather crest on the Crested Auklet…”). This reliance on vague textual features can lead to coarse reasoning and eventual misclassification by the model (e.g., wrongly identifying the bird as a “Rhinoceros Auklet” based on an imprecise assessment of “robust body” and “horn-like projection”). In contrast, as shown in Tables 14 and 16, MPO establishes a synergistic interaction between modalities. It augments a concise textual instruction with a direct visual reference, eliminating linguistic ambiguity and instructing the model to ground its reasoning in concrete visual evidence (e.g., “compare the size and shape of the bill with those in the reference image”). This multimodal grounding shifts the model’s focus, leading to more specific and visually-supported reasoning (e.g., “The bill is relatively short and thick, with a slight curve at the tip. This matches the description of the ‘crested auklet’ in the reference image…”). In addition to this, importantly, on examples from the Auklet subtask where SEE’s optimized prompt fails, MPO’s multimodal approach successfully fixes 66.6% of the misclassifications, demonstrating that multimodal cues can effectively resolve errors that are inherently difficult for text-only optimization.
@@
| Task Example: CUB(Grosbeak) Classification | |
| Optimized Prompt (SEE) (Score: 96.7) | Identify the species of the bird in the given image using the options provided. Please respond with the bird’s name exactly as it is listed (e.g., ‘rose breasted grosbeak’). Make sure to write the name precisely as it appears in the choices, without any extra punctuation or modifications. Pay attention to distinctive features such as color patterns, markings, and size to accurately identify the bird. If you find that the bird in the image does not match any of the listed options, please state ‘not identifiable’ rather than guessing. Ensure that your answer is selected only from the choices provided and refrain from inferring details about the species. After determining the bird, verify that your response is among the listed options before you submit it. |
| Image Query |
Ground Truth: Evening Grosbeak
Key Features: • White wing patches • Yellow body color |
| Response | rose breasted grosbeak |
| Task Example: CUB (Grosbeak) Classification | |
| Optimized Prompt (MPO) (Score: 100.0) |
Classify the bird in the target image by comparing it with the hybrid reference image of grosbeaks. Follow these refined steps for accurate classification.
1. Identify the Grosbeak Group: Refer to the hybrid image that displays the Rose Breasted Grosbeak, Pine Grosbeak, Blue Grosbeak, and Evening Grosbeak. Familiarize yourself with the specific traits of each species, including color patterns and markings.
2. Analyze Visual Features: Focus on these critical features of the target bird:
– Dominant Color and Markings: Note the primary color and any distinctive patterns, such as throat colors or wing designs.
– Beak Characteristics: Compare the shape and size of the beak with those in the reference image, as these can vary significantly among species.
– Body Size Comparison: Assess the body size of the target bird relative to the reference birds, ensuring accurate size comparisons.
3. Feature Prioritization:
– Prioritize color patterns first, as they are often the most telling feature.
– If colors are similar, evaluate beak shape and size next.
– Finally, consider body size. If the target bird does not closely match any reference species, provide the name of the closest match or indicate ‘unknown’, based on the following criteria:
– Closeness is determined by the degree of similarity across all analyzed features, with color being the primary factor, followed by beak shape and size.
After analyzing these features, provide the name of the bird species that most closely matches the visual characteristics observed in the target image, supported by specific observations from the hybrid reference image.
| Image Query |
Ground Truth: Evening Grosbeak
Key Features: • White wing patches • Yellow body color |
| Response | The bird in the target image has a predominantly yellow body with black wings and a white patch on its chest. Its beak is short and conical, suitable for seed-eating. Comparing these features with the hybrid reference image of grosbeaks, the bird most closely resembles the Evening Grosbeak. The yellow coloration, black wings, and white chest patch are characteristic of this species. Therefore, the bird in the target image is classified as an Evening Grosbeak. |
| Task Example: CUB(Auklet) Classification | |
| Optimized Prompt (SEE) (Score: 50.0) | Classify the content of the target image based on the provided choices of bird species: [‘parakeet auklet’, ‘rhinoceros auklet’, ‘crested auklet’, ‘least auklet’]. Comprehensive Guidelines for Classification: 1. Physical Characteristics: - Parakeet Auklet: Look for a small bird (approx. 23 cm), with a striking blue bill and a notable yellow eyebrow stripe. Its compact body and vibrant colors are key identifiers. - Rhinoceros Auklet: This medium-sized bird (about 35 cm) has a stout body and a distinctive horn-like projection on its bill during breeding season. It typically weighs around 700 grams, so note its robust build. - Crested Auklet: Identify this species by its long, curved crest of feathers and dark plumage, roughly 30 cm in size. Weighing about 300 grams, it is often found in colonies, so observe any social behaviors. - Least Auklet: The smallest of the auklets (around 20 cm), it features short wings and a stubby bill, weighing about 130 grams. Its petite size and subtle coloration are distinguishing traits. 2. Habitat and Behavior: Focus on typical habitats such as coastal regions and rocky cliffs. Note behaviors like diving and feeding. For instance, the Rhinoceros Auklet may exhibit aggressive diving, while the Crested Auklet shows unique social interactions. 3. Visual Cues: Pay attention to specific visual markers such as bill shape, color, and the feather crest on the Crested Auklet. Consistently assess the color of the feet and overall body size. 4. Common Confusions: Be aware that the Rhinoceros Auklet appears more robust than the slender Least Auklet. Additionally, plumage colors may vary under different lighting conditions. 5. Cross-Verification Process: Use a methodical approach to verify classification. For instance, check if dark plumage corresponds with the bill shape and size of the suspected species, and look for unique features. 6. Key Features to Focus On: Highlight aspects such as beak shape, feather patterns, and eye color during evaluation. 7. Narrow Down Options: If the image shows features that could apply to multiple species, focus on the most probable candidates based on observed visual cues and behaviors. 8. Confidence Level: Indicate your confidence in the classification as “certain,” “likely,” “uncertain,” or “doubtful.” Justify your confidence by detailing which specific features influenced your assessment. 9. Step-by-Step Checklist for Classification: Use this systematic checklist: Observe the size and body structure; Note the bill’s color and shape; Examine plumage patterns and unique features (e.g., crests); Consider the observed habitat and behaviors; Cross-reference traits with the characteristics of all species. 10. Hypothetical Example Scenarios: Provide examples of images and a breakdown of classifications according to the guidelines, illustrating effective application of features and observations. Please provide the identified species name and your confidence level in the classification. |
| Image Query |
Ground Truth: Crested Auklet
Key Features: • Dark plumage • Curved crest of feathers • Orange beak |
| Response | Based on the provided image, let’s analyze the bird step-by-step using the comprehensive guidelines: 1. Physical Characteristics: The bird has a robust body, which aligns with the Rhinoceros Auklet being described as medium-sized (about 35 cm) and having a stout body. The bill is orange with a black tip, which is characteristic of the Rhinoceros Auklet. There is a noticeable horn-like projection on the bill, which is a key feature of the Rhinoceros Auklet during the breeding season. 2. Habitat and Behavior: The bird is perched on a rocky cliff, which is consistent with the Rhinoceros Auklet’s habitat preference. The posture and behavior do not provide additional specific clues but do not contradict the identification either. 3. Visual Cues: The bill shape and color match the description of the Rhinoceros Auklet. The overall body structure and size are consistent with the Rhinoceros Auklet. 4. Common Confusions: The Least Auklet is much smaller and has a shorter bill, which does not match the image. The Crested Auklet has a long, curved crest, which is not present here. The Parakeet Auklet is smaller and has a distinct yellow eyebrow stripe, which is not seen here. 5. Cross-Verification Process: The combination of the robust body, orange bill with a black tip, and the horn-like projection strongly supports the identification of the Rhinoceros Auklet. 6. Key Features to Focus On: The bill shape and color are critical features for this identification. 7. Narrow Down Options: The image clearly matches the Rhinoceros Auklet based on the physical characteristics and visual cues. 8. Confidence Level: Given the clear match with the Rhinoceros Auklet’s physical characteristics and the absence of any conflicting features, I am confident in the classification. Conclusion: The identified species is the Rhinoceros Auklet, and my confidence level in this classification is certain. |
| Task Example: CUB(Auklet) Classification | |
| Optimized Prompt (MPO) (Score: 80.3) |
Analyze the provided hybrid image of auklet species and classify the bird in the target image into one of the following categories: [‘parakeet auklet’, ‘rhinoceros auklet’, ‘crested auklet’, ‘least auklet’]. Focus on the following critical visual features:
1. Bill Shape: Prioritize the size and shape of the bill, comparing it directly with the examples in the hybrid image. Be specific about curvature and thickness.
2. Plumage Color: Observe the overall body color and any variations, noting how these compare to the reference images.
3. Body Patterns: Identify unique markings or patterns on the body, ensuring to highlight any significant differences that may aid in classification.
Use the hybrid image as a guide for your comparison and classification. Clearly mention any distinguishing traits that support your decision, and ensure to prioritize bill shape in cases of ambiguity.
| Image Query |
Ground Truth: Crested Auklet
Key Features: • Dark plumage • Curved crest of feathers • Orange beak |
| Response | The bird in the target image can be classified as a “crested auklet” based on the following observations: |
| 1. Bill Shape: The bill is relatively short and thick, with a slight curve at the tip. This matches the description of the “crested auklet” in the reference image, which has a similar bill shape but with a more pronounced curve. The “rhinoceros auklet” also has a thick bill but with a more pronounced curve, while the “parakeet auklet” and “least auklet” have thinner bills. | |
| 2. Plumage Color: The overall body color is dark, almost black, with some lighter areas around the neck and underparts. This is consistent with the “crested auklet” in the reference image. The “rhinoceros auklet” has a similar dark plumage, but the “parakeet auklet” and “least auklet” have lighter plumage. | |
| 3. Body Patterns: The most distinctive feature of the “crested auklet” is the presence of a crest on its head, which is clearly visible in the target image. The “parakeet auklet” and “least auklet” do not have crests, and the “rhinoceros auklet” has a shorter crest compared to the “crested auklet”. | |
| In conclusion, the combination of the short and thick bill, dark plumage, and the presence of a crest on the head strongly supports the classification of the bird in the target image as a “crested auklet”. | |